OpenStudy (anonymous):

Help-I need to graph sin and cos type 2 y=-2sin(3/4x) 0 | 0 pi/2 | 2 pi | 0 3pi/2| -2 2pi | 0 For some reason I am wrong? Can someone help me.

OpenStudy (xapproachesinfinity):

is this a calculus quesion

OpenStudy (anonymous):

its trig

OpenStudy (xapproachesinfinity):

actually your function is undefined in 0 why do you have f(0)=0?

OpenStudy (xapproachesinfinity):

you will need more than just trig to graph this does your question actually ask to graph? perhaps you should try graphing it by calculator to graph this you need advanced techniques by hand

OpenStudy (xapproachesinfinity):

eh is the inside part (-3/4)x or -3/4x it is actually ambiguous

OpenStudy (xapproachesinfinity):

well! i believe if it's 3/4x you can draw it

OpenStudy (anonymous):

it's positive.

OpenStudy (xapproachesinfinity):

ah darn i'm sleepy i mean (3/4)x

OpenStudy (xapproachesinfinity):

so now it makes sense that 0 gives 0 but pi/2 how did you get 2?

OpenStudy (dumbcow):

hmm did my last post not help? its the period again

OpenStudy (xapproachesinfinity):

the period here is 8pi/3

OpenStudy (anonymous):

@dumbcow it did work. I tried msging to tell you Thank you! the thing is i have four more to do different problems, and if i mess up it starts me off at four again. @xapproachesinfinity here is an example of my previous question

OpenStudy (dumbcow):

ok good ... now it should work for every problem. just plug in the different amplitude and period here you have A = -2 B = 3/4 ----> period = 2pi/(3/4) = 8pi/3 0 | 0 period/4 | A period/2 | 0 (3/4)period | -A period | 0

OpenStudy (xapproachesinfinity):

follow what @dumbcow said

OpenStudy (anonymous):

ok so how did you guys get 8pi/3???

OpenStudy (xapproachesinfinity):

3/4 x=2pi then x=2pi/3/4=8pi/3

OpenStudy (xapproachesinfinity):

the period of sinx is 2pi we use that to find the period sin(3/4 x)

OpenStudy (xapproachesinfinity):

@dumbcow explained to you that if we have y=Bsin(Ax) the period is 2pi/A

OpenStudy (anonymous):

ok it's wrong i did (0,0),(2pi/3,2),(4pi/3,0),(2pi,-2),(8pi/3,0)

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

oh its because A = -2 , this means the graph is being reflected stick to the pattern but make sure to include the neg .... so (-A) = -(-2) = 2

OpenStudy (dumbcow):

in other words flip the 2 and -2 around, the graph should go down first then up

OpenStudy (anonymous):

ok so then (0,0), 2pi/3,-2), (4pi/3,0),(2pi,2),(8pi/3,0)